Reports 1-1 of 1 Clear search Modify search
VIS (BS)
takafumi.ushiba - 10:43 Thursday 05 February 2026 (36291) Print this report
TF measurement result for BS local control noise budget

I checked the measured TF of BS for the noise budget.
Following local controls have a large coherence to DARM, so precise projection should be performed.
The other TF have low coherence, so the noise contribution to the current DARM should be small.

TM_OLDAMP_P: fig1
TM_OLDAMP_Y: fig2

Images attached to this report
Comments to this report:
takaaki.yokozawa - 11:24 Thursday 05 February 2026 (36292) Print this report
I performed the noise projection manually (10 - 100 Hz).
Images attached to this comment
Search Help
×

Warning

×